A Nurse must mix 15L of a 10% solution of a drug with some 60% solution to obtain a 20% mixture. How many liters of the 60% solution will be needed?
Answer by stanbon(75887)

A Nurse must mix 15L of a 10% solution of a drug with some 60% solution to obtain a 20% mixture. How many liters of the 60% solution will be needed?
--------------------
Equation:
active + active = active
0.10*15 + 0.60x = 0.20(x+15)
Multiply thru by 100 to get:
10*15 + 60x = 20x + 20*15
40x = 10*15
x = 3.75 Liters (amt. of 60% solution needed in the mixture)
=====================
Cheers,
Stan H.